If the work function for a metal is 1.8 eV, what would be the stopping potential for light having a wavelength of 400 nm? (b) What would be the maximum velocity of the emitted photoelectrons at the metal’s surface? 


1
Expert's answer
2021-12-22T14:13:08-0500

a)

"U=\frac Ae=1.8~V,"

b)

"v=\sqrt{\frac{2hc-2A\lambda}{\lambda m}}=678~\frac{km}s."